AME: haracteristic length of the animals Animal Great white shark 15 m/swater 9x1 20 m/s Eagle Sea lion Dragonfly Dirt warm 0.001 m/s 0.05 slycerin air 1.2x 10-5 - Order the animals by the level of turbulence that they generate (from large to low). Justify your anwer5/100 polutsl . What species will experience a larger pressure drng and which ones will experience a Define DRAG and give a clear eaxplanation of the difference between presure drag and viscous dreg. Hep your explanstion by means of drawings as well [10/100 points
Answer #1

1. Turbulence is directly proportional to Reynold's number, so higher the Reynold's number, higher would be the turbulence.

ul. Re =

Order : From Highest Turbulence to lowest .

a.Eagle

Re = \frac{20*1}{1.2 * 10^-5} = 1.66667 * 10^6

b.Dragonfly

Re = \frac{7*0.1}{1.2 * 10^-5} = 5.83333 * 10^4

c.Great White Shark

Re = \frac{15*2}{9 * 10^-4} = 3.33333 * 10^4

d. Sea Lion

Re = \frac{5*3}{9 * 10^-4} = 1.66667 * 10^4

e. Dirt Warm

Re = \frac{0.001*0.05}{1} = 5 * 10^-5

2. Sea Lion will experience a greater Pressure Drag because the pressure drag is related to the surface area and since characteristic length of Sea Lion is maximum among these, Sea Lion would experience a greater pressure drag.

Eagle will experience a greater frictional drag because frictional drag scales up with the Reynold's number.

3. Drag is a force acting opposite to the relative motion of any object moving with respect to a surrounding fluid. Drag forces arise from pressure differences over the body (due to its shape) and viscous forces along the surface (in the boundary layer). Pressure Drag arises from the eddying that are set up in the fluid by the passage of the body. VIscous drag comes from the friction between the fluid and the surface over which it is flowing.